# Large Language Models Can Automatically Engineer Features for Few-Shot Tabular Learning

### v1

@@ 1 Introduction
We evaluate FeatLLM on 13 different tabular datasets in low-shot regimes, showing its strong and robust performance. We show that LLMs proficiently utilize both prior knowledge and the knowledge from the data, extracting high-quality features. Our framework outperforms contemporary few-shot learning baselines across various settings. The code is released via GitHub link at https://github.com/Sungwon-Han/FeatLLM.
@@ 3.2 Inferring Class Likelihood via Rules
To introduce randomness in rule generation for ensemble, a sufficiently high temperature is set for LLM inference. We also randomize the order of few-shot demonstrations based on the observation that the order can impact the query outcome of LLM (Lu et al., 2022) [^2] Analysis on the rule diversity can be found in Appendix E. To utilize the diversity in prediction for more robust predictions, we adopt bagging to select a subset of features or instances for each trial, which gets particularly useful for larger number of training samples or features. The proposed ensemble approach offers two main advantages. First, even if the LLM generates incorrect rules in some trials, the correct rules from other trials can compensate, enhancing the overall robustness of the framework. This ties into the LLM’s self-consistency (Wang et al., 2022), as rules commonly inferred across multiple trials are more likely to be accurate. Second, the ensemble method addresses the limitation of LLM’s prompt size. When the tabular data exceed the allowable input prompt size, bagging helps reduce it, ensuring the model’s robust performance. While a single rule set might not capture the relationships of all features, the ensemble of various weak learners created across multiple trials can nullify the impact of incomplete feature/instance coverage in individual trials.
Table 1: Evaluation results, showing AUC across 11 datasets with attributes fewer than 100, are presented. Best performances are bolded, and our framework’s performances, when second-best, are underlined. Results with more training shots (e.g., 32, 64) and an additional baseline (e.g., RandomForest) are provided in Appendix L.
@@ 4.1 Performance Evaluation
FeatLLM uses GPT-3.5 as its LLM backbone, yet is designed to be agnostic to the choice of LLMs (see Appendix I for results with different backbones). The temperature for the LLM inference is set to 0.5 and the top-p value is set to the default value of 1 in API. We set the number of ensembles and the number of rules for extracting to 20 and 10 respectively. Details on hyper-parameter impacts are in Figure 6 and Appendix B. We use the Adam optimizer of a learning rate 0.01 for the linear model, training for 200 epochs. We employ $k$-fold cross-validation for optimal epoch selection. In replicating baselines, GPT-3.5 is utilized for in-context learning, whereas T0 (Sanh et al., 2022) is employed in the TabLLM, following the original setting. Note that TabLLM restricts the use of LLM to only those with publicly available checkpoints; therefore, GPT-3.5 is not accessible in this context. Regarding traditional machine learning methods, such as LogReg, XGBoost, and RandomForest, we determined their parameters using Grid-search combined with $k$-fold cross validation. The value of $k$ is set either as 2 or 4, guaranteeing that the training set includes at least one example of each class. For more detailed information on implementations, please refer to Appendix C.
@@ 4.2 Analysis & Discussion
Figure 5 visualizes the performance changes due to spurious correlations. Compared to baselines, FeatLLM showed the least performance degradation from spurious correlations. This is likely because our framework forces the consideration of the relationship between the task and features based on prior knowledge when extracting rules, preventing the generation of rules based on irrelevant columns and thereby yielding robust results. Indeed, we observe that rules based on noisy columns accounted for only 1-2% of the total. Meanwhile, we also observe that FeatLLM can also learn spurious correlations and misunderstand causal relationships between task and features if columns from closely related data are randomly attached (see Appendix J).
@@ Appendix C Baseline Implementation Details
Appendix D Quality of Features Generated by FeatLLM
We conducted a simple experiment to investigate the informativeness of rule-based features generated by FeatLLM in solving practical tasks. Specifically, we measured the average AUROC between the newly generated features and the target labels, and then computed the ratio of useful features with an AUROC higher than 0.5 for each dataset. Table 10 below summarizes the results. It confirms that the majority of the created rules are relevant to the target attribute and provide informative insights for addressing the target task (refer to the column “Ratio of Useful Rules”). Moreover, during the process of tuning the linear model with data, rules that have little relevance to the data (i.e., rules with learned weights below the threshold) are automatically filtered out (refer to the column “Ratio of Useful Rules (after tuning)”). Note that the threshold was set at 0.1, taking into consideration the overall histogram of weights.
Table 10: Evaluation of feature quality across datasets based on AUROC.
Data Ratio of useful rules Ratio of useful rules (after tuning)
Adult 0.971 0.980
Bank 0.627 0.958
Blood 0.889 0.975
Car 0.712 0.849
Communities 0.767 0.889
Credit-g 0.643 0.802
Diabetes 0.974 0.972
Heart 0.800 0.924
Myocardial 0.690 0.744
Cultivars 0.775 0.927
NHANES 0.600 0.832
Sequence-type 0.896 0.985
Solution-mix 0.917 1.000
Average 0.789 0.911
Appendix E Diversity of Features Generated by FeatLLM
@@
Appendix F Prior Knowledge vs. Training Samples.
@@
Appendix G Error Rates of Automated Parsing via LLM
@@
Appendix H Handling missing data.
@@
Appendix I Varying LLM backbones
@@
Appendix J Additional Experiments on Effect of Adding Spurious Correlations
@@
Appendix K Qualitative Analysis
@@
Appendix L Full Results

### v2

@@ 1 Introduction
We evaluate FeatLLM on 13 different tabular datasets in low-shot regimes, showing its strong and robust performance. We show that LLMs proficiently utilize both prior knowledge and the knowledge from the data, extracting high-quality features. Our framework outperforms contemporary few-shot learning baselines across various settings. The code is released via anonymized GitHub link at https://github.com/Sungwon-Han/FeatLLM.
@@ 3.2 Inferring Class Likelihood via Rules
To introduce randomness in rule generation for ensemble, a sufficiently high temperature is set for LLM inference. We also randomize the order of few-shot demonstrations based on the observation that the order can impact the query outcome of LLM (Lu et al., 2022) [^2] Analysis on the rule diversity can be found in Appendix D. To utilize the diversity in prediction for more robust predictions, we adopt bagging to select a subset of features or instances for each trial, which gets particularly useful for larger number of training samples or features. The proposed ensemble approach offers two main advantages. First, even if the LLM generates incorrect rules in some trials, the correct rules from other trials can compensate, enhancing the overall robustness of the framework. This ties into the LLM’s self-consistency (Wang et al., 2022), as rules commonly inferred across multiple trials are more likely to be accurate. Second, the ensemble method addresses the limitation of LLM’s prompt size. When the tabular data exceed the allowable input prompt size, bagging helps reduce it, ensuring the model’s robust performance. While a single rule set might not capture the relationships of all features, the ensemble of various weak learners created across multiple trials can nullify the impact of incomplete feature/instance coverage in individual trials.
Table 1: Evaluation results, showing AUC across 11 datasets with attributes fewer than 100, are presented. Best performances are bolded, and our framework’s performances, when second-best, are underlined. Results with more training shots (e.g., 32, 64) and an additional baseline (e.g., RandomForest) are provided in Appendix K.
@@ 4.1 Performance Evaluation
FeatLLM uses GPT-3.5 as its LLM backbone, yet is designed to be agnostic to the choice of LLMs (see Appendix H for results with different backbones). The temperature for the LLM inference is set to 0.5 and the top-p value is set to the default value of 1 in API. We set the number of ensembles and the number of rules for extracting to 20 and 10 respectively. Details on hyper-parameter impacts are in Figure 6 and Appendix B. We use the Adam optimizer of a learning rate 0.01 for the linear model, training for 200 epochs. We employ $k$-fold cross-validation for optimal epoch selection. In replicating baselines, GPT-3.5 is utilized for in-context learning, whereas T0 (Sanh et al., 2022) is employed in the TabLLM, following the original setting. Note that TabLLM restricts the use of LLM to only those with publicly available checkpoints; therefore, GPT-3.5 is not accessible in this context. Regarding traditional machine learning methods, such as LogReg, XGBoost, and RandomForest, we determined their parameters using Grid-search combined with $k$-fold cross validation. The value of $k$ is set either as 2 or 4, guaranteeing that the training set includes at least one example of each class. For more detailed information on implementations, please refer to Appendix C.
@@ 4.2 Analysis & Discussion
Quality of features generated by FeatLLM.
We conduct a simple experiment to investigate the informativeness of rule-based features generated by FeatLLM in solving practical tasks. Specifically, we measure the average AUROC between the newly generated features and the target labels, and then compute the ratio of useful features with an AUROC higher than 0.5 for each dataset. Table 6 below summarizes the results. It confirms that the majority of the created rules are relevant to the target attribute and provide informative insights for addressing the target task (refer to the column “Before tuning”). Moreover, during the process of tuning the linear model with data, rules that have little relevance to the data (i.e., rules with learned weights below the threshold) are automatically filtered out (refer to the column “After tuning”). Note that the threshold was set at 0.1, taking into consideration the overall histogram of weights.
Table 6: Evaluation of feature quality before and after linear model tuning. Ratio of newly-generated features with an AUROC higher than 0.5 is reported across datasets.
Method Ratio of useful rules
Before tuning After tuning
Adult 0.971 0.980
Bank 0.627 0.958
Blood 0.889 0.975
Car 0.712 0.849
Communities 0.767 0.889
Credit-g 0.643 0.802
Diabetes 0.974 0.972
Heart 0.800 0.924
Myocardial 0.690 0.744
Cultivars 0.775 0.927
NHANES 0.600 0.832
Sequence-type 0.896 0.985
Solution-mix 0.917 1.000
Average 0.789 0.911
@@
Figure 5 visualizes the performance changes due to spurious correlations. Compared to baselines, FeatLLM showed the least performance degradation from spurious correlations. This is likely because our framework forces the consideration of the relationship between the task and features based on prior knowledge when extracting rules, preventing the generation of rules based on irrelevant columns and thereby yielding robust results. Indeed, we observe that rules based on noisy columns accounted for only 1-2% of the total. Meanwhile, we also observe that FeatLLM can also learn spurious correlations and misunderstand causal relationships between task and features if columns from closely related data are randomly attached (see Appendix I).
@@ 6 Broader Impact
Acknowledgement
We sincerely thank Scott Yak and Yihe Dong for their valuable feedback and discussion on this work.
@@ Appendix C Baseline Implementation Details
Appendix D Diversity of Features Generated by FeatLLM
@@
Appendix E Prior Knowledge vs. Training Samples.
@@
Appendix F Error Rates of Automated Parsing via LLM
@@
Appendix G Handling missing data.
@@
Appendix H Varying LLM backbones
@@
Appendix I Additional Experiments on Effect of Adding Spurious Correlations
@@
Appendix J Qualitative Analysis
@@
Appendix K Full Results